Acetone, the solvent in nail polish remover, has a density of 0.791 g/mL. what is the volume, in milliliters, of 20.0 g of acetone?


1
Expert's answer
2020-10-23T09:15:20-0400

"m=d\times V"

"V=\frac{m}{d}=\frac{20.0 g}{0.791 \frac{g}{ml}}=25.3 ml"